macmac and joshua have $500 together. if macmac will give joshua $120 they will then have equal amounts. how much has each originally
Answer by stanbon(75887)

macmac and joshua have $500 together. if macmac will give joshua $120 they will then have equal amounts. how much has each originally
----
Equations:
m + j = 500
m-120 = j+120
-----
Modify::
m + j = 500
m-j = 240
-----
Add and solve for "m"::
2m = 740
m = $370 (macmac's original amount)
----
j = 500-m = 500 - 370 = $130 (joshua's original amount)
------------
Cheers,
Stan H.
--------------